OpenStudy (anonymous):

The distance from Carrie's house to the park was 3/4 of a mile. She walked 2/3 of a mile before stopping for a drink. How far had she walked? A) 1/3 mile B) 1/2 mile C) 1/4 mile D) 5/7 mile

OpenStudy (anonymous):

@alakara_kibo @kelliegirl33

OpenStudy (anonymous):

2/3 of 3/4 = 2/3 * 3/4 = 6/12 = 1/2 mile

OpenStudy (anonymous):

thats correct